We have a solution that contains 0.30 M Zn+2 ion and 0.50 M Mn+2 ion. If you add NaOH to make the pH of the solution 10, what concentrations (M) of Zinc and Manganese will still be in solution? What percentage of the first metal was removed when the second metal starts to fall?

Using the equations if nessecary:

ΔG° = -nFE°

ln K = (nE°) / 0.0257

E = E° - (0.0257 / n) ln Q

Charge (coulombs) = Amps x time (seconds)
Answer #1

The ionizations of Zn(OH)2 and Mn(OH)2 are as below.

Zn(OH)2 (s) <========> Zn2+ (aq) + 2 OH- (aq)                      Ksp = 3*10-17

Mn(OH)2 (s) <========> Mn2+ (aq) + 2 OH- (aq)                    Ksp = 2*10-13

The pH of the solution is 10.

It is known that

pH + pOH = 14

======> 10 + pOH = 14

======> pOH = 14 – 10 = 4

Further,

pOH = -log [OH-]

where [OH-] determines the molar concentration of OH- in solution.

Therefore,

-log [OH-] = 4

======> [OH-] = antilog (-4) M

======> [OH-] = 1*10-4 M.

Plug [OH-] in the Ksp expressions and determine the concentrations of Zn2+ and Mn2+ in solution when the pH of the solution is 10.

Ksp [Zn(OH)2] = [Zn2+][OH-]2

======> 3*10-17 = [Zn2+]*(1*10-4)2 (ignore units here since Ksp is dimensionless)

======> [Zn2+] = (3*10-17)/(1*10-4)2 = 3*10-9

The concentration of Zn2+ ion remaining in solution at pH = 10 is 3*10-9 M (ans).

Ksp [Mn(OH)2] = [Mn2+][OH-]2

======> 2*10-13 = [Mn2+]*(1*10-4)2 (ignore units here since Ksp is dimensionless)

======> [Mn2+] = (2*10-13)/(1*10-4) = 2*10-5

The concentration of Mn2+ ion remaining in solution at pH = 10 is 2*10-5 M (ans).

Since Zn(OH)2 has a lower Ksp than Mn(OH)2, hence, Zn(OH)2 precipitates first. The question requires to determine [Zn2+] when Mn(OH)2 starts to precipitate.

Determine [OH-] when Mn(OH)2 starts to precipitate. Plug in [Mn2+] = 0.50 M in the Ksp expression and obtain [OH-].

Ksp = [Mn2+][OH-]2

======> 2*10-13 = (0.50)*[OH-]2

======> [OH-]2 = (2*10-13)/(0.50) = 4*10-13

======> [OH-] = 6.324*10-7

Put [OH-] = 6.324*10-7 M in the Ksp expression for Zn(OH)2 and obtain [Zn2+] when Mn(OH)2 begins to precipitate.

3*10-17 = [Zn2+][OH-]2

======> 3*10-17 = [Zn2+]*(6.324*10-7)2

======> [Zn2+] = (3*10-17)/(6.324*10-7)2 = 7.50*10-5

The concentration of free Zn2+ in the solution = 7.50*10-5 M.

The concentration of Zn2+ removed = (0.30 – 7.50*10-5) M = 0.299925 M.

Percentage of Zn2+ ion removed = (0.299925 M)/(0.30 M)*100

= 99.975%

≈ 99.97% (ans).